OpenStudy (anonymous):

Find a power series representation for the function. f(x) = 1/(9 + x) sum n=0 to infinity ____________________?

OpenStudy (anonymous):

rewrite it as \[\frac{1}{9}\left(\frac{1}{1+\frac{x}{9}}\right)\]

OpenStudy (anonymous):

then use \[\frac{1}{1-x}=\sum_{n=0}^{\infty}x^n\] replacing \(x\) by \(-\frac{x}{9}\)

OpenStudy (anonymous):

wrong:S